OpenStudy (anonymous):

Simplify: 8x3 – 2x2 + x + 3x3 A. 10x3 B. 11x3– 2x2+ x C. 11x6– 2x2+ x D. 8x3– 2x2+ x + 3x3 (the expression cannot be simplified) @ganeshie8 @ luigi0210

OpenStudy (anonymous):

will medal and fan

OpenStudy (anonymous):

When looking at \[ 8x^3-2x^2+x+3x^3\], you want to try to combine like terms. What like terms do you see?

OpenStudy (anonymous):

8x3 and 2x3 and 3x3 @onemillion6

OpenStudy (anonymous):

8x^3 and 3x^3 are like terms, yes! (But -2x^2 is not because it has a different power on the x. You want the exponents to match.) Now, we can rewrite what is given above as \[ 8x^3+3x^3-2x^2+x\] (just switching the order) You combine 8x^3+3x^3 but adding the coefficients together. So what is 8x^3+3x^3?

OpenStudy (anonymous):

would it be 10^3+3^3

OpenStudy (anonymous):

Not quite. One way to think about 8x^3+3x^3 is: \[ (x^3+x^3+x^3+x^3+x^3+x^3+x^3+x^3)+(x^3+x^3+x^3)\] There are 8 x^3's in the group inside the first parentheses and 3 in the second group. That is what 8x^3+3x^3. Looking at the equation above, how many x^3's are there in total?

OpenStudy (anonymous):

8^3+3^3?

OpenStudy (anonymous):

Ah. I think I get what you're trying to do. When you combine like terms, the x's don't just go away. So the sum you're looking for is going to be something like: (a number)x^3 What if I write it like this: \[ 8x^3+3x^3 = x^3(8+3) = ? \] Can you complete the equation?

OpenStudy (anonymous):

11^3?

OpenStudy (anonymous):

11x^3! Don't leave out the x. But, correct :) So your new expression is \[ 11x^3 -2x^2 + x \] Do you see that in your answer choices?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

hooray, then you'er done :)